# thanks: Corresponding authors:

### v1

@@ I Introduction
The repressor-operator model of Jacob and Monod provided not only a successful conceptual vision for gene expression writ large, but also served as the basis of mathematical models of transcription based upon the precepts of statistical mechanics [21, 22, 23]. These models provided a quantitative description of a variety of different regulatory contexts in which the strengths of binding sites, the repressor copy number and DNA loop length were altered, illustrating how genetic circuits could be tuned directly and quantitatively [24, 25, 26, 27, 28]. Interestingly, these pioneering studies became a jumping off point for the construction of a number of synthetic variants that when combined with fluorescent reporters made it possible to watch synthetic switches and oscillators in real time in single cells [29, 30, 31].
In addition to the seminal discoveries of the existence of gene circuits themselves, a parallel set of discoveries unfolded which added a second layer of regulatory control to the original repressor-operator model and its subsequent generalizations and elaborations. Specifically, the mystery of induction required another insight into biological feedback and control. Enzymatic adaptation, the idea that somehow enzymes that were latent would become active in the presence of the right substrate [6], led to the discovery of allostery, a concept that Monod himself referred to as the “second secret of life” [32]. In the context of gene regulation, this idea implies that transcription factors themselves are subject to control through the binding of effector molecules that alter their activity [33, 34, 35, 36, 37, 38, 39, 40]. Writ large, these insights now fall under the general heading of allosteric transitions, a phenomenon in which proteins of all types undergo conformational changes that alter their activity. This idea applies broadly to ion channels, enzymes, the respiratory protein hemoglobin, membrane receptors mediating chemotaxis and quorum sensing, and of course, to the main subject of our paper, transcription factors [40].
@@
The production rate in Eqn. 2, and throughout this work, is modeled using a thermodynamic framework relating promoter occupancy to output [21, 27, 23, 46, 47, 48]. We note that often instead of adopting the full thermodynamic model to treat promoter occupancy, it is considered convenient to use Hill functions as an approximation to describe the probability of promoter binding [42, 43, 44, 45]. The auto-activation example will serve as our first foray into the problem of induction of genetic circuits by effector molecules as well as an opportunity to bring some critical scrutiny to the use of Hill functions to describe the physics of occupancy.
@@
Typically, the exploration of the stability behavior of these circuits is based upon varying theoretically accessible parameters such as dissociation constants ($K_{d}$), transcription rates ($r_{i}$) and degradation rates ($\gamma$) as featured in Eqn. 2 and in Fig. 3(B-D), without reference to how such parameters are themselves controlled by living cells [40]. In fact, often the underlying response is dictated by the presence and absence of experimentally accessible effector molecules that alter the balance between inactive and active forms of key regulatory molecules such as transcription factors as shown in Fig. 3(A). Thus, while all of the tunable parameters in Fig. 3 make it possible to systematically tune the level of gene expression, some of these parameters are more conveniently accessible to the experimentalist and to the cell itself as it rapidly tunes its behavior in response to stimuli. Our goal is to use explicit statistical mechanical models of the induction phenomenon to explore the behavior of genetic circuits as a function of the presence or absence of effectors.
@@ II The Statistical Mechanics of Induction
We now undertake a systematic analysis of a number of different regulatory circuits from the point of view of allosteric regulation, building upon earlier work in which the biological parameters are tuned by hand rather than through the action of effectors [29, 42, 43, 49, 50]. In particular, we argue that often the number of active transcription factors $TF_{\text{act}}$ is given by
@@
where $TF_{\text{tot}}$ is the total number of transcription factors and $p_{\text{act}}(c)$ is the probability that the transcription factor is active as a function of the effector concentration $c$. We will show that the fraction of transcription factors that are active can be given by the Monod-Wyman-Changeux (MWC) model, which can be used to compute $p_{\text{act}}(c)$ using statistical mechanics [33, 34, 35, 36, 37, 38, 39, 40]. Note that although we invoke the MWC model to describe allostery, we could just as well use the KNF model or even the phenomenological Hill functions to capture the role of the effector [40].
In the MWC model, we consider an inactive state and an active state of the transcription factor with energy difference $\varepsilon=\varepsilon_{i}-\varepsilon_{a}$. The states and weights for such an allosteric transcription factor with two binding sites is shown in Fig. 4. Appealing to these states and weights, the probability of a transcription factor being active is then of the form
@@
Figure 4: States and weights for an allosteric transcription factor with an effector that can bind at two sites on the molecule. The sum of the thermodynamic weights for the active and inactive conformations are shown at the bottom.
@@
Figure 5: Activity of a transcription factor as a function of its effector concentration $c$. Unless stated otherwise, the parameters used throughout this work are: $K_{A}=140\,\mu M$, $K_{I}=530\,nM$, $\varepsilon=4.5\,k_{B}T$. (A) Probability of active transcription factor as a function of effector concentration, defined by Eqn. 4. The half maximal effective concentration EC50, defined as the effector concentration $c^{*}$ such that $p_{\text{act}}(c^{*})=(p_{\text{act}}^{max}+p_{\text{act}}^{min})/2$, is plotted in purple. (B) The effective dissociation constant $K_{d}^{\text{eff}}$ (dimensionless with respect to $K_{d}$) as a function of effector concentration. Saturation of $p_{\text{act}}(c)$ corresponds to minimal $K_{d}^{\text{eff}}$, and leakiness of $p_{\text{act}}(c)$ corresponds to maximal $K_{d}^{\text{eff}}$.
Writing $K_{d}^{\text{eff}}=K_{d}/p_{\text{act}}(c)$ in gene regulatory network motifs is useful and informative in many ways. First, traditionally, theoretical models of the behavior of network motifs are studied by tuning $K_{d}^{\text{eff}}$, while experiments generally tune the effector concentration $c$. Incorporating $p_{\text{act}}$ bridges theory and experiments by providing experimentally accessible “knobs” to control the behavior of the genetic circuit of interest. Second, since $p_{\text{act}}(c)$ is highly nonlinear in $c$, model variables might react to varying $c$ differently than varying $K_{d}^{\text{eff}}$. Third, $p_{\text{act}}(c)$ constrains the range of accessible $K_{d}^{\text{eff}}$ values. When discussing input-output curves, there are a variety of summary parameters that help us understand their character qualitatively. For example, in the case where effector binding renders the protein inactive, the leakiness is the amount of activity at saturating concentrations of effector, namely, $p_{\text{act}}(\infty)$. Similarly, the maximal activity known as the saturation occurs in the zero-effector limit, namely, $p_{\text{act}}(0)$. These important summary variables can be calculated as
@@
where $\bar{K}_{c}=K_{A}/K_{I}$. With the parameters used in Fig. 5(A), $p_{\text{act}}^{\min}$ and $p_{\text{act}}^{\max}$ are separated by about three orders of magnitude, and thus $K_{d}^{\text{eff}}$ can also only be tuned across three orders of magnitude, as shown in Fig. 5(B). The restricted $K_{d}^{\text{eff}}$ range can have important consequences. For example, consider a bistable system with a stability curve such as that shown in Fig. 6. The $x$-axis is $K_{d}^{\text{eff}}$ and the $y$-axis tracks the steady state concentration of some protein $A$. The range of $K_{d}^{\text{eff}}$ resulting in bistability might not be fully accessible, depending on the baseline $K_{d}$ value.
@@
Figure 6: Allosteric tuning restricts the range of accessible $K_{d}^{\text{eff}}$. The dark blue curve shows the steady-state values of activator concentration $A$, with filled circles indicating stable fixed points and open circles indicating unstable ones. The light blue and light green shaded regions represent two example ranges of accessible $K_{d}^{\text{eff}}=K_{d}/p_{\text{act}}(c)$ values, obtained by varying the effector concentration $c$ for two proteins with different DNA binding affinities (and thus different baseline $K_{d}$ values). For the blue region, $K_{d}=3.2\times 10^{-2}$ $\mu$M $=10^{-1.5}$ $\mu$M; for the green region, $K_{d}=1$ $\mu$M. The parameters for the auto-activation stability curve are: $\gamma=1/$min, $r_{0}=0.1$ $\mu$M///min, $r_{1}=1$ $\mu$M///min, $r_{2}=20$ $\mu$M///min, $\omega=100$. If $K_{d}$ itself were tunable, the full positive $K_{d}^{\text{eff}}$ axis would be accessible. However, tuning the biologically relevant parameter $c$ imposes constraints on the range of achievable $K_{d}^{\text{eff}}$. The light blue range does not intersect the bistable region, while the light green range does, illustrating that bistability may not be fully accessible due to the functional dependence of $K_{d}^{\text{eff}}$ on $p_{\text{act}}(c)$.
Figure 7: Mutual repression is ubiquitous in cellular decision making. (A) The bacteriophage lambda switch that mediates the phage decision to become a lysogen in the bacterium or engage in lysis through mutual repression of cI and cro (and some autoactivation of cI as well). (B) In hematopoietic development, mutual repression between different genes have been suggested to ensure the switch-like adoption of alternate cellular fates. Diagram adapted from [54] and [55]. (C) Mutual repression in the early fruit fly gene network has been associated with the emergence of discrete domains of gene expression. Diagram adapted from [56] and [57].
@@ III Bistability in Genetic Circuits
Different gene networks serve different biological functions. Among the most important classes of networks are those that yield bistability. Bistability refers to the situation in which, for a given set of parameters, a system can exist in one of two possible stable steady states. Such a feature is biologically significant. For example, often the expression level of one protein determines a cell’s fate. To obtain cells with different functions, there might be some cells with high concentrations of the protein of interest and some cells with low concentrations, requiring a bistable system regulating the protein in question. Broad computational surveys have shown that switch-like, bistable behavior can emerge from a wide range of simple biochemical network architectures [58]. Here, we analyze the two simplest and most ubiquitous gene circuits that produce bistability  [44], considered now through the new lens of how effector molecules modulate the dissociation constants of regulatory proteins binding to DNA. This is in contrast to the conventional setting in which the tuning strategy is simply to modulate the binding parameters within thermodynamic models [42, 43, 44, 49], rather than those parameters being naturally tuned through the action of signaling molecules.
@@ III.1 The auto-activation regulatory motif
Auto-activation circuits, in which a gene product enhances its own transcription, are among the simplest genetic regulatory motifs and are capable of generating bistable behavior  [42]. Such motifs have been studied extensively in both synthetic and natural biological systems. In vitro synthetic networks have demonstrated robust switching between high and low expression states under controlled biochemical conditions [59], and positive feedback loops have been implicated in natural processes like cell differentiation, where they convert graded input signals into binary gene expression responses [60]. Motivated by these observations, we now turn to a theoretical analysis of the auto-activation switch. The architecture we study is shown schematically in Fig. 2, where a transcription factor activates its own production, forming a feedback loop.
The states included in the states and weights diagram of Fig. 2 reveal that an essential element of achieving bistability is the cooperative binding of more than one activator on the regulatory region of the gene’s promoter. Given these states, weights, and rates, we can now write the kinetic equation governing the dynamics of the auto-activation system as
@@
Figure 8: Plot of production and degradation rates for an auto-activation switch as a function of activator concentration. This figure illustrates the competition between the production and degradation terms for a system with rate constants $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, $\bar{r}_{2}=20$, and cooperativity $\omega=10$ at effector concentration $c=25\;\mu$M. Intersections of the curve denote stable (filled nodes at low concentration $\bar{A}_{\text{low}}$ and high concentration $\bar{A}_{\text{high}}$) and unstable (unfilled node $\bar{A}_{\text{unstable}}$) fixed points. The vectors denote a phase portrait representing the direction and magnitude of change in activator concentration as a function of activator concentration itself.
@@
Figure 10: Plot of stable state evolution, exhibiting hysteresis under different trends for effector concentration. The blue curve plots how the stable state evolves with increasing effector concentration from an initial high concentration of activator. The orange dashed curve plots the change in stable activator concentration starting from an initially low level as effector concentration decreases. Comparing to the bifurcation diagram (grey) previously shown in Fig. 9, the history of steady-state evolution determines the threshold concentration of effector at which the system switches state, highlighting hysteretic behavior.
Note that the threshold at which the system switches from one stable state to another differs when increasing and decreasing the effector concentration. If the system initially contains a high concentration of activator before tuning, a higher concentration of effector is necessary to switch to the low activator state than is required when decreasing effector concentration for a system with initially low activator concentration. Considering Fig. 9 again, the higher threshold corresponds to the maximum effector concentration at which the system is bistable, and the lower threshold to the minimum concentration at which the system is bistable. Fig.  10 illustrates this phenomenon of hysteresis more explicitly, overlaying the previously discussed bifurcation diagram, and showing how the threshold at which the blue stable equilibrium trajectory switches from high to low activator concentration differs from the threshold for the orange trajectory tracing the switch in reverse.
@@
Figure 11: Range of effector concentrations for which the system exhibits bistability. The red shaded region indicates bistability. Outside this region, the system is monostable. Unless otherwise specified, parameters are fixed at: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, and $\bar{r}_{2}=20$. (A) Bistable concentration range as a function of cooperativity $\omega$, varied over the interval $\omega\in[10^{-6},10^{6}]$, corresponding to interaction energies between the two activators from approximately $+14,k_{B}T$ to $-14,k_{B}T$, since $\omega=e^{-\beta\varepsilon_{\text{int}}}$. The dotted line shows the analytical lower bound on the minimal cooperativity required for bistability. (B) Bistable concentration range as a function of $\bar{r}_{0}$, sampled over the interval $\bar{r}_{0}\in[10^{-6},\bar{r}_{1}]$. (C) Bistable concentration range as a function of $\bar{r}_{1}$, sampled over the interval $\bar{r}_{1}\in[\bar{r}_{0},\bar{r}_{2}]$. (D) Bistable concentration range as a function of $\bar{r}_{2}$, sampled over the interval $\bar{r}_{2}\in[\bar{r}_{1},10^{6}]$. The rate parameters are varied under the constraint of the auto-activation condition $\bar{r}_{0}\leq\bar{r}_{1}\leq\bar{r}_{2}$, which ensures that the production rate increases with the number of bound activators.

### v2

@@ I Introduction
The repressor-operator model of Jacob and Monod provided not only a successful conceptual vision for gene expression writ large, but also served as the basis of mathematical models of transcription based upon the precepts of statistical mechanics [1, 88, 10]. These models provided a quantitative description of a variety of different regulatory contexts in which the strengths of binding sites, the repressor copy number and DNA loop length were altered, illustrating how genetic circuits could be tuned directly and quantitatively [69, 65, 103, 102, 46, 81, 99, 100]. Interestingly, these pioneering studies became a jumping off point for the construction of a number of synthetic variants that when combined with fluorescent reporters made it possible to watch synthetic switches and oscillators in real time in single cells [27, 23, 70].
In addition to the seminal discoveries of the existence of gene circuits themselves, a parallel set of discoveries unfolded which added a second layer of regulatory control to the original repressor-operator model and its subsequent generalizations and elaborations. Specifically, the mystery of induction required another insight into biological feedback and control. Enzymatic adaptation, the idea that somehow enzymes that were latent would become active in the presence of the right substrate [67], led to the discovery of allostery, a concept that Monod himself referred to as the “second secret of life” [98]. In the context of gene regulation, this idea implies that transcription factors themselves are subject to control through the binding of effector molecules that alter their activity [28, 61, 62, 56, 57, 13, 29, 74, 68, 20]. Writ large, these insights now fall under the general heading of allosteric transitions, a phenomenon in which proteins of all types undergo conformational changes that alter their activity. This idea applies broadly to ion channels, enzymes, the respiratory protein hemoglobin, membrane receptors mediating chemotaxis and quorum sensing, and of course, to the main subject of our paper, transcription factors [74].
@@
The production rate in Eqn. 2, and throughout this work, is modeled using a thermodynamic framework relating promoter occupancy to output [39, 40, 102, 10, 7, 6, 90]. We note that often instead of adopting the full thermodynamic model to treat promoter occupancy, it is convenient to use Hill functions as an approximation to describe the probability of promoter binding [14, 2, 17, 25, 8, 77, 96, 95, 104]. The auto-activation example will serve as our first foray into the problem of induction of genetic circuits by effector molecules as well as an opportunity to bring some critical scrutiny to the use of Hill functions to describe the physics of occupancy.
@@
Typically, the exploration of the stability behavior of these circuits is based upon varying theoretically accessible parameters such as dissociation constants ($K_{d}$), transcription rates ($r_{i}$) and degradation rates ($\gamma$) as featured in Eqn. 2 and in Fig. 3(B-D), without reference to how such parameters are themselves controlled by living cells [74]. In fact, often the underlying response is dictated by the presence and absence of experimentally accessible effector molecules that alter the balance between inactive and active forms of key regulatory molecules such as transcription factors, as shown in Fig. 3(A). Thus, while all of the tunable parameters in Fig. 3 make it possible to systematically tune the level of gene expression, some of these parameters are more conveniently accessible to the experimentalist and to the cell itself as it rapidly tunes its behavior in response to stimuli. Our goal is to use explicit statistical mechanical models of the induction phenomenon to explore the behavior of genetic circuits as a function of the presence or absence of effectors.
@@ II The Statistical Mechanics of Induction
We now undertake a systematic analysis of a number of different regulatory circuits from an allosteric perspective, building upon earlier work in which biological parameters are tuned by hand rather than by effectors [27, 14, 2, 73, 92, 70]. Although some previous studies also vary inducer concentration, they typically either address different regulatory contexts than ours, map inducer concentrations to protein activity without modeling the allosteric transition explicitly, or consider different types of dynamical systems analyses than we do [105, 67, 42, 70, 81, 99, 101, 104, 47, 58, 19, 18, 77].
In particular, we argue that often the number of active transcription factors $TF_{\text{act}}$ is given by
@@
where $TF_{\text{tot}}$ is the total number of transcription factors and $p_{\text{act}}(c)$ is the probability that the transcription factor is active as a function of the effector concentration $c$. For example, effector binding can render a protein inactive, decreasing $p_{\text{act}}$ as effector concentration $c$ increases. We will show that the fraction of transcription factors that are active can be given by the Monod-Wyman-Changeux (MWC) model, which can be used to compute $p_{\text{act}}(c)$ using statistical mechanics [68, 28, 61, 62, 56, 57, 13, 29, 74]. Note that while we invoke the MWC model to describe allostery, we could just as well use the KNF model or even the phenomenological Hill functions to capture the role of the effector [74, 68]. The KNF model, for example, would describe effector binding as sequential and local. In this case, transcription factor structure would change at the subunit level, rather than through a single coordinated change in the protein’s quartenary conformation as described by the MWC interpretation [73, 22].
Figure 4: States and weights for an allosteric transcription factor with an effector that can bind at two sites on the protein. Effectors can bind to both the active and inactive forms of the transcription factor with different dissociation constants $K_{A}$ and $K_{I}$, which determine whether the effector stabilizes the protein more strongly in its active or inactive configuration. The sum of the thermodynamic weights for the active and inactive conformations are shown at the bottom.
In the MWC model, we consider an inactive state and an active state of the transcription factor with energy difference $\varepsilon=\varepsilon_{i}-\varepsilon_{a}$. In this setting, an effector can bind to a transcription factor in both its active and inactive forms with different dissociation coefficients. The values of these constants then determine whether increasing effector concentration stabilizes the active or inactive form of the transcription factor. The states and weights for such an allosteric transcription factor with two binding sites is shown in Fig. 4. Appealing to these states and weights, the probability of a transcription factor being active is then of the form
@@
Figure 5: Activity of a transcription factor as a function of effector concentration $c$. Parameters used throughout the paper, unless otherwise stated, are: $K_{A}=140\,\mu M$, $K_{I}=530\,nM$, $\varepsilon=4.5\,k_{B}T$. (A) Probability of active transcription factor as a function of effector concentration, defined by Eqn. 4. The half maximal effective concentration EC50, defined as the effector concentration $c^{*}$ such that $p_{\text{act}}(c^{*})=(p_{\text{act}}^{max}+p_{\text{act}}^{min})/2$, is plotted in purple. (B) The effective dissociation constant $K_{d}^{\text{eff}}$ (dimensionless with respect to $K_{d}$) as a function of effector concentration. Saturation of $p_{\text{act}}(c)$ corresponds to minimal $K_{d}^{\text{eff}}$, and leakiness of $p_{\text{act}}(c)$ corresponds to maximal $K_{d}^{\text{eff}}$.
Writing $K_{d}^{\text{eff}}=K_{d}/p_{\text{act}}(c)$ in gene regulatory network motifs is useful and informative in many ways. First, traditionally, theoretical models of the behavior of network motifs are studied by tuning $K_{d}^{\text{eff}}$, while experiments generally tune the effector concentration $c$. Incorporating $p_{\text{act}}(c)$ bridges theory and experiments by providing experimentally accessible “knobs” to control the behavior of the genetic circuit of interest. Second, since $p_{\text{act}}(c)$ is highly nonlinear in $c$, other model variables might react differently to varying $c$ than to varying $K_{d}^{\text{eff}}$. Third, $p_{\text{act}}(c)$ constrains the range of accessible $K_{d}^{\text{eff}}$ values. When discussing input-output curves, there are a variety of summary parameters that help us understand their character qualitatively. For example, in the case where effector binding renders the protein inactive, the leakiness is the amount of activity at saturating concentrations of effector, namely, $p_{\text{act}}(\infty)$. Similarly, the maximal activity known as the saturation occurs in the zero-effector limit, namely, $p_{\text{act}}(0)$. These important summary variables can be calculated as
@@
where $\bar{K}_{c}=K_{A}/K_{I}$. With the parameters used in Fig. 5(A), $p_{\text{act}}^{\min}$ and $p_{\text{act}}^{\max}$ are separated by about three orders of magnitude, and thus $K_{d}^{\text{eff}}$ is also only tunable across three orders of magnitude, as shown in Fig. 5(B). The restricted $K_{d}^{\text{eff}}$ range can have important consequences. For example, consider a bistable system with a stability curve such as that shown in Fig. 6, where the $x$-axis is $K_{d}^{\text{eff}}$ and the $y$-axis tracks the steady state concentration of some protein $A$. Tuning the parameter $c$ imposes constraints on the range of achievable $K_{d}^{\text{eff}}$ for a given $K_{d}$. The light blue range shown does not intersect the bistable regime, while the light green range does, illustrating that bistability may not be fully accessible due to the functional dependence of $K_{d}^{\text{eff}}$ on $p_{\text{act}}(c)$. This restriction on $K_{d}$ follows a tradition of relating thermodynamic binding interactions to experimentally measured biochemical parameters [82].
@@
[…]
Figure 7: Mutual repression is ubiquitous in cellular decision making. (A) The bacteriophage lambda switch that mediates the phage decision to become a lysogen in the bacterium or engage in lysis through mutual repression of cI and cro (and some auto-activation of cI as well). (B) In hematopoietic development, mutual repression between different genes have been suggested to ensure the switch-like adoption of alternate cellular fates. Diagram adapted from [48] and [15]. (C) Mutual repression in the early fruit fly gene network has been associated with the emergence of discrete domains of gene expression. Diagram adapted from [11] and [43].
@@ III Bistability in Genetic Circuits
Different gene networks serve different biological functions. Among the most important classes of networks are those that yield bistability. Bistability refers to the situation in which, for a given set of parameters, a system can exist in one of two possible stable steady states. Such a feature is biologically significant. For example, often the expression level of one protein determines a cell’s fate. To obtain cells with different functions, there might be some cells with high concentrations of the protein of interest and some cells with low concentrations, requiring a bistable system regulating the protein in question. Broad computational surveys have shown that switch-like, bistable behavior can emerge from a wide range of simple biochemical network architectures [91]. Here, we analyze the two simplest and most ubiquitous gene circuits that produce bistability [17], auto-activation and mutual repression, now through the new lens of how effector molecules modulate the dissociation constants of regulatory proteins binding to DNA. Although there are examples in the literature that also use this knob in other regulatory or modeling contexts (as noted in Section II), the approach considered here offers a contrast to the conventional setting often employed in studying auto-activation and mutual repression. In such cases, the tuning strategy is typically to modulate the binding parameters within thermodynamic models [14, 2, 17, 73], rather than those parameters being naturally tuned through the action of signaling molecules. For simplicity, the following sections do not consider transcription factor oligomerization, either in solution or when bound to DNA, which could give rise to more complex configurations such as DNA looping. We discuss modest generalizations to this effect in Appendix G.
@@ III.1 The auto-activation regulatory motif
Auto-activation circuits, in which a gene product enhances its own transcription, are among the simplest genetic regulatory motifs and are capable of generating bistable behavior [14]. Such motifs have been studied extensively in both synthetic and natural biological systems. In vitro synthetic networks have demonstrated robust switching between high and low expression states under controlled biochemical conditions [94], and positive feedback loops have been implicated in natural processes like cell differentiation, where they convert graded input signals into binary gene expression responses [5]. The inducible auto-activation circuit has also been explored with Hill function formulations and with $p_{act}$ as an input parameter [36]. Motivated by this work, we now present our theoretical analysis of the auto-activation switch that incorporates a full thermodynamic model and explicit treatment of induction. The architecture we study is shown schematically in Fig. 2, where a transcription factor activates its own production, forming a feedback loop. Given the states, weights, and rates in Fig. 2, we can now write the kinetic equation governing the dynamics of the auto-activation system as
@@
[…]
@@
Figure 10: Plot of stable state evolution, exhibiting hysteresis under different trends for effector concentration. The blue curve plots how the stable state evolves with increasing effector concentration from an initial high concentration of activator. The orange dashed curve plots the change in stable activator concentration starting from an initially low level as effector concentration decreases. Comparing to the bifurcation diagram (gray) previously shown in Fig. 9, the history of steady-state evolution determines the threshold concentration of effector at which the system switches state, highlighting hysteretic behavior.
Note that the threshold at which the system switches from one stable state to another differs when increasing and decreasing the effector concentration. If the system initially contains a high concentration of activator before tuning, a higher concentration of effector is necessary to switch to the low activator state than is required when decreasing effector concentration for a system with initially low activator concentration. Considering Fig. 9 again, the higher threshold corresponds to the maximum effector concentration at which the system is bistable, and the lower threshold to the minimum concentration at which the system is bistable. Fig. 10 illustrates this phenomenon of hysteresis more explicitly, overlaying the previously discussed bifurcation diagram, and showing how the threshold at which the blue stable equilibrium trajectory switches from high to low activator concentration differs from the threshold for the orange trajectory tracing the switch in reverse.
@@
Figure 11: Range of effector concentrations for which the system exhibits bistability. The red shaded region indicates bistability. Outside this region, the system is monostable. Unless otherwise specified, parameters are fixed at: $\omega=7.5$, $\bar{r}_{0}=0.1$, $\bar{r}_{1}=1$, and $\bar{r}_{2}=20$. (A) Bistable concentration range as a function of cooperativity $\omega$, varied over the interval $\omega\in[10^{-6},10^{6}]$, corresponding to interaction energies between the two activators from approximately $14,k_{B}T$ to $14,k_{B}T$, since $\omega=e^{-\beta\varepsilon_{\text{int}}}$. The dotted line shows the analytical lower bound on the minimal cooperativity required for bistability. (B) Bistable concentration range as a function of $\bar{r}_{0}$, sampled over the interval $\bar{r}_{0}\in[10^{-6},\bar{r}_{1}]$. (C) Bistable concentration range as a function of $\bar{r}_{1}$, sampled over the interval $\bar{r}_{1}\in[\bar{r}_{0},\bar{r}_{2}]$. (D) Bistable concentration range as a function of $\bar{r}_{2}$, sampled over the interval $\bar{r}_{2}\in[\bar{r}_{1},10^{6}]$. The rate parameters are varied under the constraint of the auto-activation condition $\bar{r}_{0}\leq\bar{r}_{1}\leq\bar{r}_{2}$, which ensures that the production rate increases with the number of bound activators.